You drive 9 miles east then 9 miles north


1
Expert's answer
2022-02-20T15:47:46-0500

Answer

You drive 9 miles east then 9 miles north

Totoal dispalcement is given by

"D=\sqrt{9^2+9^2}\\=1.414*9=12.70miles."